A bacterium moves with a speed of 3.5  across a petri dish with a radius of 8.4 cm. How long does it take the bacterium to traverse the petri dish? (Hint: assume it travels through the centre of the dish.)



1
Expert's answer
2021-03-09T15:27:27-0500

The total distance the bacterium has to travell is "d = 2\cdot 8.4cm = 16.8cm" (the diameter of the dish). I assume that the speed of the bacterium is "v = 3.5cm\/min". Then the time it takes the bacterium to traverse the petri dish is:


"t = \dfrac{d}{v} = \dfrac{16.8cm}{3.5cm/min} = 4.8min"

Answer. 4.8 minutes.